OpenStudy (anonymous):

What is the prime factorization of 72? WILL GIVE MEDAL AND FAN! 1x2x4x9 22*x32* 22*x33* 23*x32* *This number is a exponent

OpenStudy (nykoreyan):

2^3 x 3^2

OpenStudy (alexandervonhumboldt2):

72/2=36 36/2=18 18/2=9 9/3=3 3/3=1 1*3*3*2*2*2=1*9*2*4=1*2*4*9

OpenStudy (alexandervonhumboldt2):

or 2^3*3^2

OpenStudy (anonymous):

What nykoreyan said?